# Automorphisms of Plane Curves defined from Chebychev polynomials

### v1

@@ Abstract.
In this paper, we study the geometry and automorphism groups of the algebraic curves $\mathcal{C}_{d}$ defined by the equation $y^{d}=\varphi_{d}(x)$ over a field $k$ with $\operatorname{char}(k)\nmid 2d$, where $\varphi_{d}(x)$ is the Chebyshev polynomial of degree $d$. We classify the total inflection points of $\mathcal{C}_{d}$, correcting and extending previous work on this. Additionally, we determine the automorphism groups of $\mathcal{C}_{d}$ in several cases, namely for $d=4$, and for any $d$ such that $12d=q+$ 1 or $14d=q+$ 1 for an arbitrary power $q$ of the prime $p=\operatorname{char}(k)$. As an application, we use our results to show that certain maximal curves (over finite fields) of the same genus are not isomorphic.
@@ 1. Introduction and preliminaries
Our main results concern the classification of total inflection points of $\mathcal{C}_{d}$ and a detailed description of $\operatorname{Aut}({\mathcal{C}}_{d})$ in several key cases. Specifically, we prove the following:
@@
    The total inflection points of $\mathcal{C}_{d}$ are precisely the $d$ points in $\mathcal{Z}(y)\cap\mathcal{C}_{d}$, except in the case where $\operatorname{char}(k)=p\geq 3$ and $d=\frac{p^{m}+1}{2}$ for some $m\geq 1$. In this exceptional case there are $3d$ total inflection points. This result corrects and extends [1, Theorem 6.3] (see Proposition 2.1).
@@
    If $12d=q+$ 1 for some power $q$ of $p$, which is the situation described in (I), then $\mathcal{C}_{d}$ is isomorphic to a Fermat curve of degree $d$ and its automorphism group is $(\mathbb{Z}/d\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z})\rtimes S_{3}$ (see Propositions 3.1 and 3.3).
@@
    If $14d=q+$ 1 for some power $q$ of $p$, then the automorphism group of $\mathcal{C}_{d}$ is $\mathbb{Z}/d\mathbb{Z}\rtimes S_{3}$ (see Theorem 3.17 and Proposition 3.18).
@@
    We show that for many integers $m$ and $n$ and finite fields ${\mathbb{F}}_{q^{2}}$, the maximal curves defined by $y^{n}=\varphi_{m}(x)$ and $y^{n}=x^{m}+1$ are not isomorphic despite having the same genus. This result provides a corrected proof of [1, Corollary 6.4] and also generalizes it (see Propositions 4.1, 4.2, and 4.3).
The organization of this paper is as follows. In Section 2, we analyze the total inflection points of $\mathcal{C}_{d}$ and establish the correction to [1]. Section 3 is dedicated to the classification of the automorphism groups in the cases $12d=q+$ 1 and $14d=q+$ 1, respectively. Finally, Section 4 discusses applications to maximal curves over finite fields.
@@ 2. Total inflection points
Since automorphisms of any such smooth plane curve ${\mathcal{C}}$ are given by linear maps on ${\mathbb{P}}^{2}$, they act as permutations on the set of total inflection points of ${\mathcal{C}}$. We will compute these points in the case of ${\mathcal{C}}_{d}$, assuming $d\geq 3$ and $\operatorname{char}(k)\nmid 2d$. The result will help us in our ultimate goal, namely obtaining information on $\operatorname{Aut}({\mathcal{C}}_{d})$. The computation extends the proof of [1, Theorem 6.3], where the special case $p:=\operatorname{char}(k)>2$ and $\overline{-1}\in\langle\,\overline{p}\,\rangle\subseteq(\mathbb{Z}/2d\mathbb{Z})^{\times}$ is considered. Moreover, our computation corrects some claims on p. 17 of [1]: for example, the reasoning presented there leads to the erroneous conclusion that ${\mathcal{C}}_{d}$ has either $d$ or $2d$ total inflection points, contradicting their assertion in [1, Theorem 6.3].
@@
We consider all possibilities for a line $\ell\subset{\mathbb{P}}^{2}$ with #$\#(\ell\cap{\mathcal{C}}_{d})=1$.
@@
(4) What remains, is the case $\ell\colon y=ax+bz$ where $a\neq 0$. This gives #$\#(\ell\cap{\mathcal{C}}_{d})=1$ precisely in the situation
@@
We claim that $c_{1}\neq 0$. Indeed, the assumption $c_{1}=0$ implies that $\varphi_{d}(x)=(ax+b)^{d}-c_{2}^{d}$. Now observe that the two maps $t\mapsto\varphi_{d}(t)$ and $t\mapsto(at+b)^{d}-c_{2}^{d}$ are clearly distinct (the first has only $\infty$ as a ramification point with index $d$ while the other is totally ramified over both $\infty$ and $-c_{2}^{d}$). As a consequence,
@@
    with $i$ any primitive $4$-the root of unity. Write $d=3+qn$ with $q$ a power of $p$ and $\gcd(n,p)=1$. Comparing coefficients of $t^{qn-q+2}$ results in ±$\pm 6n\equiv 0\bmod p$, a contradiction. Hence this case cannot occur.
@@ 3. Automorphism groups
over a field of characteristic $p>0$. If $12d=q+$ 1 with $q$ a power of $p$, then $\mathcal{C}_{d}$ is isomorphic over $\mathbb{F}_{q^{2}}$ to the Fermat curve
@@
in case $12d=q+$ 1. Equivalently, we claim that for $d=(q+1)/2$, in the polynomial ring ${\mathbb{F}}_{p}[u]$ the equality $(u-1)^{d}\varphi_{d}(\frac{2u+2}{u-1})=2u^{d}+2$ holds. Indeed, write $u=s^{2}$. In $\mathbb{F}_{p}(s)\supset{\mathbb{F}}_{p}(u)$ one observes
@@
For $12d=q+$ 1, it is known that the curve $y^{d}=\varphi_{d}(x)$ is maximal over $\mathbb{F}_{q^{2}}$. Thus, by [2, Theorem 1.1], the curve is isomorphic to the maximal Fermat curve of degree $d=(q+1)/2$. Here we provided a direct proof of this result.
@@
Consider the curve $\mathcal{C}_{d}$ given by $y^{d}=\varphi_{d}(x)$ defined over a field of characteristic $p>0$. If $12d=q+$ 1 with $q$ a power of $p$, then the automorphism group $G$ of the curve $\mathcal{C}_{d}$ has order $6d^{2}$ and has a normal abelian subgroup $H\cong\mathbb{Z}/d\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z}$ of order $d^{2}$ such that $G/H\cong S_{3}.$
@@
±$\pm 2dab^{d-1}=0.$
This implies that $a=0$. Now comparing the terms of degree $d-2$ shows $-d=-db^{2}$, leading to $b=\pm 1$.
@@
Let $S=\{\alpha,-\alpha,1/\alpha,-1/\alpha\}$ be a subset of some field $k$, and assume #$\#S=4$ (equivalently, $\alpha^{4}\neq 1$). Then the group
@@
Restriction to $S$ identifies $G_{S}$ with a subgroup of $S_{4}$. Under this identification, the group $\langle\sigma,\tau\rangle$ generated by $\sigma$ and $\tau$ corresponds to the normal subgroup of $S_{4}$ generated by (1 2)(3 4)1234(1\,2)(3\,4)( 1 2 ) ( 3 4 ) and (1 3)(2 4)1324(1\,3)(2\,4)( 1 3 ) ( 2 4 ), of order $4$. If $G_{S}$ is strictly larger than $\langle\sigma,\tau\rangle$, then either it contains a $2$-cycle, or a $3$-cycle. We now analyse these two possibilities.
If $G_{S}$ contains a $2$-cycle, after possibly multiplying by an element of $\langle\sigma,\tau\rangle$ we can assume it fixes $\alpha$ and one more element of $S$. In particular, considered as an $f\in\operatorname{Aut}({\mathbb{P}}^{1})$ we have $f(0)\neq 0$, which means $f$ has the form $f\colon x\mapsto\frac{ax+1}{cx+d}$. Moreover $f$ having order $2$ implies $d=-a$. Three possibilities remain:
[…]
The fixpoints show $f(x)=\alpha^{2}/x$ and therefore $-1/\alpha=f(1/\alpha)=\alpha^{3}$. This means that one is in case (1) of the lemma.
[…]
@@
[…]
@@
Note that any $3$-cycle has two fixpoints in ${\mathbb{P}}^{1}$. For $f$, one is $\alpha$ and we claim that the other one is not $\infty$. Indeed, $f(\infty)=\infty$ means that $f\colon x\mapsto ax+b$ for certain $a\neq 0,b$. Considering how $f$ acts on ±$\pm\alpha$ and $1/\alpha$ leads to the conclusion $(\alpha^{2}-1)^{2}=0$, contradicting the assumption that #$\#S=4$.
@@
Let $k$ be an algebraically closed field $k$ of characteristic 2,5,7$\neq$ 2$\neq$ 5$\neq$ 7≠ 2 , ≠ 5 , ≠ 7. The automorphism group $G$ of the curve $y^{4}=\varphi_{4}(x)$ over $k$ has order $16$ and contains a normal subgroup $N\cong\mathbb{Z}/4\mathbb{Z}$ of order $4$ such that
@@
Since for $d=4$ we have $12d=p+$ 1 for $p=7$, Proposition 2.1 and our assumption that $\operatorname{char}(k)\neq 7$ imply that ${\mathcal{C}}_{4}$ has precisely $4$ total inflection points. Hence by Lemma 3.4 an exact sequence
@@
Hence, writing $N$ the normal subgroup of $G=\operatorname{Aut}({\mathcal{C}}_{4})$ given above, by Lemma 3.9 and the observation $2\lambda^{2}\neq 0$, the group $G/N$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ except when $2\lambda^{2}=\pm 6$ or $x^{4}-2\lambda^{2}x^{2}+1$ divides $x^{8}+14x^{4}+1$. The first exception implies $836=4\lambda^{4}=$ 8, which does not occur by the assumption on $\operatorname{char}(k)$. Since $x^{8}+14x^{4}+1$ results in the remainder $2040\lambda^{2}x^{2}-$ 20 when divided by $x^{4}-2\lambda^{2}x^{2}+1$ and $\operatorname{char}(k)\neq 2$, the remaining exception is only possible for $\operatorname{char}(k)=5$.
@@
Assume $14d=p^{r}+$ 1 and $d>4$. Denote by $G$ the automorphism group of the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ (in characteristic $p$), and let the subgroup $N<G$ be given by
@@
Since $14d-$ 1 and $12d-$ 1 cannot both be powers of $p$, Proposition 2.1 shows that ${\mathcal{C}}_{d}$ contains precisely $d$ total inflection points. Hence Lemma 3.4 implies that $N$ is normal in $G$. Consider the element $\sigma\colon x\mapsto-x$ in $G/N$. From Lemma 3.12, if there exists an element $\eta\colon x\mapsto\frac{1}{bx}$ in $G/N$ for some constant $b$, then either $p=3$ or $d\equiv 4$ or $d\equiv\frac{1}{2}\pmod{p}$. We claim that $p\neq 3$. Indeed the condition $14d=3^{r}+$ 1 implies that $d$ is odd, contradicting Lemma 3.12.
Since $14d=p^{r}+$ 1, we have $d\not\equiv\frac{1}{2}\pmod{p}$. Hence $d\equiv 4\pmod{p}$, implying $16\equiv 4d\equiv 1\pmod{p}$, so $p=5$. However, this is impossible since the condition $14d=p^{r}+$ 1 implies $p\equiv-1\pmod{4}$.
@@
Consider the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ defined over a field $k$ of characteristic $p$, with $14d=p^{r}+$ 1 for some $r>0$. Then ${\mathcal{C}}_{d}$ admits the automorphism of order $3$ given by
@@
Write $q=p^{r}$, then $14d=q+$ 1 and therefore $\sqrt{-1}\not\in{\mathbb{F}}_{q}$. This implies
@@
$2F\left(s^{4}+\frac{1}{s^{4}}\right)=\left(s+\frac{1}{s}\right)^{4d}+\left(s-\frac{1}{s}\right)^{4d}=2\left(s^{d}+\frac{1}{s^{d}}\right),$
where it is used that $14d=q+$ 1. As a consequence $F(x)=\varphi_{d}(x)$, completing the proof. ∎
@@
As a polynomial over $\mathbb{Q}$, the splitting field of $\varphi_{d}(x)$ equals $\mathbb{Q}(\zeta+\zeta^{-1})$ with Galois group over $\mathbb{Q}$ equal to $({\mathbb{Z}}/4d{\mathbb{Z}})^{\times}/\{\pm 1\}$. The splitting field of $\varphi_{d}(x)$ over ${\mathbb{F}}_{p}$ therefore has Galois group isomorphic to the decomposition group at $p$ in $({\mathbb{Z}}/4d{\mathbb{Z}})^{\times}/\{\pm 1\}$, which is the subgroup generated by $\bar{p}$. The equality $14d=p^{r}+$ 1 shows that this subgroup has order $r$, hence the splitting field of $\varphi_{d}(x)$ over ${\mathbb{F}}_{p}$ equals ${\mathbb{F}}_{q}$. In particular, this means that $c\in{\mathbb{F}}_{q}^{\times}$. From the proof of Proposition 3.15, as $14d=q+$ 1, one obtains
@@
for some constant $\lambda\neq 0$. Squaring both sides and substituting $x=t+\frac{1}{t}$ (note that $14d=q+$ 1, so $\left((A+B)^{2d}+(A-B)^{2d}\right)^{2}=(A+B)^{q+1}+(A-B)^{q+1}+2(A^{2}-B^{2})^{2d}=2(A^{2})^{2d}+2(B^{2})^{2d}+2(A^{2}-B^{2})^{2d}$ and also $44^{2d}=2^{q+1}=$ 4 ), gives
@@
    Suppose all three are squares, or all three are nonsquares. Then $01-6c+2-1-6c=$ 0, hence $p>3$ and $c=\frac{1}{6}$. Substituting this into the $\deg(t)=2$ equation leads to $p=7$ and $c=-1$ (and $d\equiv 2\bmod 7$). Hence one considers $\mu\colon x\mapsto(x-2)/(1-x)$ in characteristic $7$, which has order $6$ and fixpoints ±$\pm 3$ (the zeros of $\varphi_{2}(x)=x^{2}-2$). As $3$ is not a square in ${\mathbb{F}}_{q}$ for $q$ an odd power of $7$, we have $3^{2d}\equiv-3\bmod 7$ and hence here
@@
    In case only $1-2c$ is a nonsquare or is the only square, one obtains $01-6c+2+1+6c=$ 0, again a contradiction.
@@
    In case only $c$ is a nonsquare and also when only $c$ is a square, we have $01-6c-2-1-6c=$ 0 hence $p>3$ and $c=-\frac{1}{6}$. This results using the $t^{2}$-coefficient in $p=7$, so $c=1$. We have the inverse of the map $\mu$ discussed in case (a) here, resulting in the same contradiction.
@@
    If only $1+2c$ is a nonsquare, and also if only $1+2c$ is a square, then the $t$-coefficient does not give a further restriction. From the $t^{2}$-coefficient one concludes $012c^{2}+1=$ 0. This violates that $\mu\in\text{Aut}({\mathbb{P}}^{1})$, i.e., the matrix $\begin{pmatrix}1&-12c\\ c&1\end{pmatrix}$ is invertible.
@@
Suppose that the characteristic of the field $k$ is $p\geq 3$. If $14d=q+$ 1, where $q$ is a power of $p$ and $d\geq 4$, then the automorphism group $G$ of the curve
@@
satisfies $\mu r\mu r=\textit{id}$, one concludes in case $\mu\neq r,\textit{id}$ that $a=d$. Assuming that moreover $\mu\in{\mathbb{Z}}/n{\mathbb{Z}}<D_{n}$, so that $\mu s=s\mu$, one derives $b=-12c$, so
@@
Assume $\operatorname{char}(k)=p>3$, and let $d$ be an integer such that $3\nmid d$. If $14d=p^{r}+$ 1, then the automorphism group $G$ of the curve $y^{d}=\varphi_{d}(x)$ has order $6d$ and contains a normal subgroup $N$ of order $d$ such that $G/N\cong D_{3}$.

### v2

# Automorphisms of Plane Curves defined from Chebyshev polynomials

@@ Abstract.
We investigate the automorphism groups of the algebraic curves
$\mathcal{C}_{d}:y^{d}=\varphi_{d}(x),$
where $\varphi_{d}(x)$ denotes the Chebyshev polynomial of degree $d$, defined over a field $k$ with $p:=\operatorname{char}(k)\nmid 2d$. We determine the full automorphism group of $\mathcal{C}_{d}$ in all the cases considered in this paper, namely for $d=4$, and more generally when $2d=p^{r}+1$ or $4d=p^{r}+1$. For all other $d>4$, Expectation 3.19 predicts what the automorphism group should be.
As an application, we show that certain maximal curves of the same genus are not isomorphic.
@@ 1. Introduction and preliminaries
The main goal of this paper is to determine the automorphism groups of the curves ${\mathcal{C}}_{d}$ in several significant cases. As a tool in our study, we revisit the classification of the total inflection points of ${\mathcal{C}}_{d}$ and provide a transparent proof. Specifically, we prove the following:
@@
    The total inflection points of $\mathcal{C}_{d}$ are precisely the $d$ points in $\mathcal{Z}(y)\cap\mathcal{C}_{d}$, except in the case where $\operatorname{char}(k)=p\geq 3$ and $d=\frac{p^{m}+1}{2}$ for some $m\geq 1$. In this exceptional case there are $3d$ total inflection points. This result extends [1, Theorem 6.3] (see Proposition 2.1).
@@
    If $2d=q+1$ for some power $q$ of $p$, which is the situation described in (I), then $\mathcal{C}_{d}$ is isomorphic to a Fermat curve of degree $d$ and its automorphism group is $(\mathbb{Z}/d\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z})\rtimes S_{3}$ (see Propositions 3.1 and 3.3); Although this fact follows from the general results of [5], we provide here a short and elementary proof which avoids the machinery used in that work.
@@
    If $4d=q+1$ for some power $q$ of $p$, then the automorphism group of $\mathcal{C}_{d}$ is $\mathbb{Z}/d\mathbb{Z}\rtimes S_{3}$ (see Theorem 3.17 and Proposition 3.18).
@@
    We show that for many integers $m$ and $n$ and finite fields ${\mathbb{F}}_{q^{2}}$, the maximal curves defined by $y^{n}=\varphi_{m}(x)$ and $y^{n}=x^{m}+1$ are not isomorphic despite having the same genus. This result provides a refined proof of [1, Corollary 6.4] and also generalizes it (see Propositions 4.1, 4.2, and 4.3).
The organization of this paper is as follows. In Section 2, we analyze the total inflection points of $\mathcal{C}_{d}$ and clarify certain aspects of [1]. Section 3 is dedicated to the classification of the automorphism groups in the cases $2d=q+1$, $d=4$ and $4d=q+1$, respectively. Finally, Section 4 discusses applications to maximal curves over finite fields.
@@ 2. Total inflection points
Since automorphisms of any such smooth plane curve $\mathcal{C}$} are given by linear maps on ${\mathbb{P}}^{2}$, they act as permutations on the set of total inflection points of $\mathcal{C}$}. We will compute these points in the case of ${\mathcal{C}}_{d}$, assuming $d\geq 3$ and $\operatorname{char}(k)\nmid 2d$. The result will help us in our ultimate goal, namely obtaining information on $\operatorname{Aut}({\mathcal{C}}_{d})$. The computation extends the proof of [1, Theorem 6.3], where the special case $p:=\operatorname{char}(k)>2$ and $\overline{-1}\in\langle\,\overline{p}\,\rangle\subseteq(\mathbb{Z}/2d\mathbb{Z})^{\times}$ is considered. Moreover, our computations offer a refined analysis of some statements in [1, p. 17]. For instance, the reasoning there seems to suggest that $\mathcal{C}_{d}$ has either $d$ or $2d$ total inflection points, which does not align with the result in [1, Theorem 6.3].
@@
We consider all possibilities for a line $\ell\subset{\mathbb{P}}^{2}$ with $(\ell\cap{\mathcal{C}}_{d})=1$.
@@
(4) What remains, is the case $\ell\colon y=ax+bz$ where $a\neq 0$. This gives $(\ell\cap{\mathcal{C}}_{d})=1$ precisely in the situation
@@
We claim that $c_{1}\neq 0$. Indeed, the assumption $c_{1}=0$ implies that $\varphi_{d}(x)=(ax+b)^{d}-c_{2}^{d}$. Now observe that the two maps $t\mapsto\varphi_{d}(t)$ and $t\mapsto(at+b)^{d}-c_{2}^{d}$ are clearly distinct (the first has only $\infty$ as a ramification point with index $d$ while the other is totally ramified over both $\infty$ and $c_{2}^{d}$). As a consequence,
@@
    with $i$ any primitive $4$-the root of unity. Write $d=3+qn$ with $q$ a power of $p$ and $\gcd(n,p)=1$. Comparing coefficients of $t^{qn-q+2}$ results in $\pm 6n\equiv 0\bmod p$, a contradiction. Hence this case cannot occur.
@@ 3. Automorphism groups
over a field of characteristic $p>0$. If $2d=q+1$ with $q$ a power of $p$, then $\mathcal{C}_{d}$ is isomorphic over $\mathbb{F}_{q^{2}}$ to the Fermat curve
@@
in case $2d=q+1$. Equivalently, we claim that for $d=(q+1)/2$, in the polynomial ring ${\mathbb{F}}_{p}[u]$ the equality $(u-1)^{d}\varphi_{d}(\frac{2u+2}{u-1})=2u^{d}+2$ holds. Indeed, write $u=s^{2}$. In $\mathbb{F}_{p}(s)\supset{\mathbb{F}}_{p}(u)$ one observes
@@
For $2d=q+1$, it is known that the curve $y^{d}=\varphi_{d}(x)$ is maximal over $\mathbb{F}_{q^{2}}$. Thus, by [2, Theorem 1.1], the curve is isomorphic to the maximal Fermat curve of degree $d=(q+1)/2$. Here we provided a direct proof of this result.
@@
Consider the curve $\mathcal{C}_{d}$ given by $y^{d}=\varphi_{d}(x)$ defined over a field of characteristic $p>0$. If $2d=q+1$ with $q$ a power of $p$, then the automorphism group $G$ of the curve $\mathcal{C}_{d}$ has order $6d^{2}$ and has a normal abelian subgroup $H\cong\mathbb{Z}/d\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z}$ of order $d^{2}$ such that $G/H\cong S_{3}.$
@@
$\pm 2dab^{d-1}=0.$
This implies that $a=0$. Now comparing the terms of degree d2d-2 shows $d=-db^{2}$, leading to $b=\pm 1$.
@@
Let $S=\{\alpha,-\alpha,1/\alpha,-1/\alpha\}$ be a subset of some field $k$, and assume #S=4\#S=4 (equivalently, $\alpha^{4}\neq 1$). Then the group
@@
Restriction to $S$ identifies $G_{S}$ with a subgroup of $S_{4}$. Under this identification, the group $\langle\sigma,\tau\rangle$ generated by $\sigma$ and $\tau$ corresponds to the normal subgroup of $S_{4}$ generated by (1 2)(3 4)(1\,2)(3\,4) and (1 3)(2 4)(1\,3)(2\,4), of order $4$. If $G_{S}$ is strictly larger than $\langle\sigma,\tau\rangle$, then either it contains a $2$-cycle, or a $3$-cycle. We now analyse these two possibilities.
If $G_{S}$ contains a $2$-cycle, after possibly multiplying by an element of $\langle\sigma,\tau\rangle$ we can assume it fixes $\alpha$ and one more element of $S$. In particular, considered as an $f\in\operatorname{Aut}({\mathbb{P}}^{1})$ we have $f(0)\neq 0$, which means $f$ has the form $f\colon x\mapsto\frac{ax+1}{cx+d}$. Moreover $f$ having order $2$ implies d=ad=-a. Three possibilities remain:
$\text{(a)}\quad\quad\qquad\alpha\;\text{and}\;-\alpha\;\text{fixed,}\;\;1/\alpha\mapsto-1/\alpha\mapsto 1/\alpha.$
The fixpoints show $f(x)=\alpha^{2}/x$ and therefore $1/\alpha=f(1/\alpha)=\alpha^{3}$. This means that one is in case (1) of the lemma.
$\text{(b)}\quad\quad\qquad\alpha\;\text{and}\;1/\alpha\;\text{fixed,}\;\;-\alpha\mapsto-1/\alpha\mapsto-\alpha.$
@@
$\text{(c)}\quad\quad\qquad\alpha\;\text{and}\;-1/\alpha\;\text{fixed,}\;\;-\alpha\mapsto 1/\alpha\mapsto-\alpha.$
@@
Note that any $3$-cycle has two fixpoints in ${\mathbb{P}}^{1}$. For $f$, one is $\alpha$ and we claim that the other one is not $\infty$. Indeed, $f(\infty)=\infty$ means that $f\colon x\mapsto ax+b$ for certain $a\neq 0,b$. Considering how $f$ acts on $\pm\alpha$ and $1/\alpha$ leads to the conclusion $(\alpha^{2}-1)^{2}=0$, contradicting the assumption that #S=4\#S=4.
@@
Let $k$ be an algebraically closed field $k$ of characteristic $\neq 2,\neq 5,\neq 7$. The automorphism group $G$ of the curve $y^{4}=\varphi_{4}(x)$ over $k$ has order $16$ and contains a normal subgroup $N\cong\mathbb{Z}/4\mathbb{Z}$ of order $4$ such that
@@
Since for $d=4$ we have $2d=p+1$ for $p=7$, Proposition 2.1 and our assumption that $\operatorname{char}(k)\neq 7$ imply that ${\mathcal{C}}_{4}$ has precisely $4$ total inflection points. Hence by Lemma 3.4 an exact sequence
@@
Hence, writing $N$ the normal subgroup of $G=\operatorname{Aut}({\mathcal{C}}_{4})$ given above, by Lemma 3.9 and the observation $2\lambda^{2}\neq 0$, the group $G/N$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ except when $2\lambda^{2}=\pm 6$ or $x^{4}-2\lambda^{2}x^{2}+1$ divides $x^{8}+14x^{4}+1$. The first exception implies $36=4\lambda^{4}=8$, which does not occur by the assumption on $\operatorname{char}(k)$. Since $x^{8}+14x^{4}+1$ results in the remainder $40\lambda^{2}x^{2}-20$ when divided by $x^{4}-2\lambda^{2}x^{2}+1$ and $\operatorname{char}(k)\neq 2$, the remaining exception is only possible for $\operatorname{char}(k)=5$.
@@
Assume $4d=p^{r}+1$ and $d>4$. Denote by $G$ the automorphism group of the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ (in characteristic $p$), and let the subgroup $N<G$ be given by
@@
Since 4d14d-1 and 2d12d-1 cannot both be powers of $p$, Proposition 2.1 shows that ${\mathcal{C}}_{d}$ contains precisely $d$ total inflection points. Hence Lemma 3.4 implies that $N$ is normal in $G$. Consider the element $\sigma\colon x\mapsto-x$ in $G/N$. From Lemma 3.12, if there exists an element $\eta\colon x\mapsto\frac{1}{bx}$ in $G/N$ for some constant $b$, then either $p=3$ or $d\equiv 4$ or $d\equiv\frac{1}{2}\pmod{p}$. We claim that $p\neq 3$. Indeed the condition $4d=3^{r}+1$ implies that $d$ is odd, contradicting Lemma 3.12.
Since $4d=p^{r}+1$, we have $d\not\equiv\frac{1}{2}\pmod{p}$. Hence $d\equiv 4\pmod{p}$, implying $16\equiv 4d\equiv 1\pmod{p}$, so $p=5$. However, this is impossible since the condition $4d=p^{r}+1$ implies $p\equiv-1\pmod{4}$.
@@
Consider the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ defined over a field $k$ of characteristic $p$, with $4d=p^{r}+1$ for some $r>0$. Then ${\mathcal{C}}_{d}$ admits the automorphism of order $3$ given by
@@
Write $q=p^{r}$, then $4d=q+1$ and therefore $\sqrt{-1}\not\in{\mathbb{F}}_{q}$. This implies
@@
$2F\left(s^{4}+\frac{1}{s^{4}}\right)=\left(s+\frac{1}{s}\right)^{4d}+\left(s-\frac{1}{s}\right)^{4d}=2\left(s^{4d}+\frac{1}{s^{4d}}\right),$
where it is used that $4d=q+1$. As a consequence $F(x)=\varphi_{d}(x)$, completing the proof. ∎
@@
As a polynomial over $\mathbb{Q}$, the splitting field of $\varphi_{d}(x)$ equals $\mathbb{Q}(\zeta+\zeta^{-1})$ with Galois group over $\mathbb{Q}$ equal to $({\mathbb{Z}}/4d{\mathbb{Z}})^{\times}/\{\pm 1\}$. The splitting field of $\varphi_{d}(x)$ over ${\mathbb{F}}_{p}$ therefore has Galois group isomorphic to the decomposition group at $p$ in $({\mathbb{Z}}/4d{\mathbb{Z}})^{\times}/\{\pm 1\}$, which is the subgroup generated by $\bar{p}$. The equality $4d=p^{r}+1$ shows that this subgroup has order $r$, hence the splitting field of $\varphi_{d}(x)$ over ${\mathbb{F}}_{p}$ equals ${\mathbb{F}}_{q}$. In particular, this means that $c\in{\mathbb{F}}_{q}^{\times}$. From the proof of Proposition 3.15, as $4d=q+1$, one obtains
@@
for some constant $\lambda\neq 0$. Squaring both sides and substituting $x=t+\frac{1}{t}$ (note that $4d=q+1$, so $\left((A+B)^{2d}+(A-B)^{2d}\right)^{2}=(A+B)^{q+1}+(A-B)^{q+1}+2(A^{2}-B^{2})^{2d}=2(A^{2})^{2d}+2(B^{2})^{2d}+2(A^{2}-B^{2})^{2d}$ and also $4^{2d}=2^{q+1}=4$ ), gives
@@
    Suppose all three are squares, or all three are nonsquares. Then 16c+216c=01-6c+2-1-6c=0, hence $p>3$ and $c=\frac{1}{6}$. Substituting this into the $\deg(t)=2$ equation leads to $p=7$ and c=1c=-1 (and $d\equiv 2\bmod 7$). Hence one considers $\mu\colon x\mapsto(x-2)/(1-x)$ in characteristic $7$, which has order $6$ and fixpoints $\pm 3$ (the zeros of $\varphi_{2}(x)=x^{2}-2$). As $3$ is not a square in ${\mathbb{F}}_{q}$ for $q$ an odd power of $7$, we have $3^{2d}\equiv-3\bmod 7$ and hence here
@@
    In case only 12c1-2c is a nonsquare or is the only square, one obtains 16c+2+1+6c=01-6c+2+1+6c=0, again a contradiction.
@@
    In case only $c$ is a nonsquare and also when only $c$ is a square, we have 16c216c=01-6c-2-1-6c=0 hence $p>3$ and $c=-\frac{1}{6}$. This results using the $t^{2}$-coefficient in $p=7$, so $c=1$. We have the inverse of the map $\mu$ discussed in case (a) here, resulting in the same contradiction.
@@
    If only $1+2c$ is a nonsquare, and also if only $1+2c$ is a square, then the $t$-coefficient does not give a further restriction. From the $t^{2}$-coefficient one concludes $12c^{2}+1=0$. This violates that $\mu\in\text{Aut}({\mathbb{P}}^{1})$, i.e., the matrix $\begin{pmatrix}1&-12c\\ c&1\end{pmatrix}$ is invertible.
@@
Suppose that the characteristic of the field $k$ is $p\geq 3$. If $4d=q+1$, where $q$ is a power of $p$ and $d\geq 4$, then the automorphism group $G$ of the curve
@@
satisfies $\mu r\mu r=\textit{id}$, one concludes in case $\mu\neq r,\textit{id}$ that $a=d$. Assuming that moreover $\mu\in{\mathbb{Z}}/n{\mathbb{Z}}<D_{n}$, so that $\mu s=s\mu$, one derives b=12cb=-12c, so
@@
Assume $\operatorname{char}(k)=p>3$, and let $d$ be an integer such that $3\nmid d$. If $4d=p^{r}+1$, then the automorphism group $G$ of the curve $y^{d}=\varphi_{d}(x)$ has order $6d$ and contains a normal subgroup $N$ of order $d$ such that $G/N\cong D_{3}$.